OpenStudy (anonymous):

Solve the equation: 4x + 9 = -3

OpenStudy (anonymous):

subtract 9 and then divide by 4. x= -3

OpenStudy (mimi_x3):

4x + 9 = -3 4x = -3 -9 4x = -12 x = -3